# A Generative Approach for Wikipedia-Scale Visual Entity Recognition

### 04 Mar 2024 v1

@@ Abstract
Code will be released.
@@ 3.1 Problem definition
We design the code set $\mathcal{C}$ so that it has three properties which we find are important for effective ger models: i) semantically structured thanks to language, ii) discriminative and compact, and iii) unambiguous. Our algorithm to create such unambiguous, language-based and discriminative codes, called ald, is illustrated in Fig. 2 (a), described in pseudo-code in Algorithm 1 and detailed in the following.
@@ 3.3 Training
Entity-based pre-training. Common auto-regressive models such as GIT [45] or PaLI [5] are pre-trained for descriptive captioning. As shown in Tab. 5 and Fig 9, they generalize poorly to entity recognition. This is because of the task discrepancy between predicting a descriptive caption and predicting an entity name (see examples in Fig. 9). In order to condition our models better for entity recognition, we propose to collect a significant number of entity-based pretraining images, each associated with a Wikipedia entity instead of a generic caption. However, such an entity-based pretraining dataset does not exist. We create it in an automatic way, without any human supervision.
@@ 3.4 Baselines
ger-baselines: alternative code creation strategies. We compare ger-ald, i.e. the best variant of ger, with several alternatives. First, ger-atomic refers to using atomic, completely unstructured codes, i.e. each code token $c^{e}_{i}$ is randomly drawn from $\llbracket 1,V\rrbracket^{L}$ [42] . Second, we consider two alternatives using semantically structured codes: (i) ger-hkc where we embed all the entity names with a pretrained text encoder before applying hierarchical k-means clustering on the resulting embeddings [42] and (ii) ger-caption where we create a code by tokenizing the entity name with $\Phi(.)$ [12, 6]. More details on these baselines are in Sec. 6.4.
@@ 4.1 Experimental setting
Pretraining and finetuning. Unless specified otherwise, we pretrain our models on the entity-WebLI dataset, which we create considering all 6M entity names as described in Sec. 3.3. After this entity-based pretraining, the models are finetuned on OVEN training set which consists only of the “seen” entities. All implementation details are in Sec. 6 and code will be released in scenic library [7].
Preventing data leakage. We remove pretraining images from Entity-WebLI and Entity-LAION with a cosine similarity (with CLIP-L/14 visual features) above 0.95 with any of the OVEN test or val images. We chose a 0.95 conservative threshold by looking at some examples: similarity 0.95 corresponds to conceptually similar images but clearly not duplicates (see Fig. 8).
@@ 4.4.1 Semantic versus atomic codes
In Fig. 3 (and Tab. 6), we report the relative improvement of semantically-structured codes (ger-ald) compared to unstructured codes (ger-atomic). We vary: (i) pretraining data size, (ii) model capacity and (iii) label-space size. A relative improvement of $100\%$ means that the performance of ger-ald is doubled compared to ger-atomic.
@@
Entities with long entity names. In Fig. 4 (left), we report the accuracy per entity name length for both ger-ald and ger-caption finetuned models. We see that the longer the entity name, the more ger-ald improves over captioning. Longer entities tend to have more noise with key information further into the code. We also show in Fig. 4 qualitative examples of entities with long entity names (more in Fig. 12). In the left example, we see that ger-ald use the token combination [col][ob] to represent the semantic concept of colobus monkey species. The last token is used to efficiently differentiate between sub-species of colobus. This compact and discriminative way of encoding the entity allows ger-ald to successfully predict this entity whereas ger-caption fails to generate the entity tokenized name.
@@ 4.4.3 Creating codes with ald
Least frequent tokens. In Tab. 3 (left), we validate our choice of selecting the least frequent tokens by evaluating 3 alternatives: random choice, most frequent tokens and first-appearing tokens in the tokenized entity name. We see that these alternative selection strategies hurt the performance significantly. Qualitative examples in Fig. 11 show that the kept tokens are less semantic and discriminative compared to ger-ald strategy of keeping the least frequent tokens. Note that all these variants are at least as good as ger-atomic ($11.4$ top-1) which is not based on language at all.
@@ 4.5 Link with classification
A typical way of tackling visual entity recognition is by training a classifier into the number of entities [35]. This is not a viable solution for web-scale problems such as OVEN where a single fully-connected layer for a $6$M classes has an enormous parameter count of $4.6$B. In this section, we evaluate ger in cases where learning a classification model is a feasible choice (smaller number of classes). Classification can be cast in our ger framework simply by setting $L=1$ and $V=|\mathcal{E}|=$ number of classes (see Sec. 3.1), making it a special case of atomic codes with $L=1$. Since the decoder decodes a single token, it is equivalent to a multi-layer Multihead Attention Pooling (MAP) head [48, 21]. In Tab. 4, we consider two challenging classification datasets: long-tailed ImageNet-LT [24] and noisy Webvision [22]. We evaluate ger-{ald, atomic} and a classification baseline using multi-layer perceptron (MLP) on averaged-pooled patch tokens. Implementation details are in Sec 6.3.

### 21 Mar 2024 v2

@@ Abstract
Code: github.com/google-research/scenic/tree/main/scenic/projects/gerald
@@ 3.1 Problem definition
We design the code set $\mathcal{C}$ so that it has three properties which we find are important for effective ger models: i) semantically structured thanks to language, ii) discriminative and compact, and iii) unambiguous. Our algorithm to create such unambiguous, language-based and discriminative codes, called ald, is illustrated in Fig. 2 (a) and described in pseudo-code in Algorithm 1 of the Appendix.
@@ 3.3 Training
Entity-based pre-training. Common auto-regressive models such as GIT [45] or PaLI [5] are pre-trained for descriptive captioning. As shown in Tab. 5 and Fig 9 of the Appendix, they generalize poorly to entity recognition. This is because of the task discrepancy between predicting a descriptive caption and predicting an entity name. In order to condition our models better for entity recognition, we propose to collect a significant number of entity-based pretraining images, each associated with a Wikipedia entity instead of a generic caption. However, such an entity-based pretraining dataset does not exist. We create it in an automatic way, without any human supervision.
@@ 3.4 Baselines
ger-baselines: alternative code creation strategies. We compare ger-ald, i.e. the best variant of ger, with several alternatives. First, ger-atomic refers to using atomic, completely unstructured codes, i.e. each code token $c^{e}_{i}$ is randomly drawn from $\llbracket 1,V\rrbracket^{L}$ [42] . Second, we consider two alternatives using semantically structured codes: (i) ger-hkc where we embed the entity names with a pretrained text encoder before applying hierarchical k-means clustering on the resulting embeddings [42] and (ii) ger-caption where we create a code by tokenizing the entity name with $\Phi(.)$ [12, 6]. Details on the baselines are in Appendix Sec. 6.4.
@@ 4.1 Experimental setting
Pretraining and finetuning. Unless specified otherwise, we pretrain our models on the entity-WebLI dataset, which we create considering all 6M entity names as described in Sec. 3.3. After this entity-based pretraining, the models are finetuned on OVEN training set which consists only of the “seen” entities. All implementation details are in Sec. 6 in Appendix and code is released in the scenic library [7].
Preventing data leakage. We remove pretraining images from Entity-WebLI and Entity-LAION with a cosine similarity (with CLIP-L/14 visual features) above 0.95 with any of the OVEN test or val images. We chose a 0.95 conservative threshold by looking at some examples: similarity 0.95 corresponds to conceptually similar images but clearly not duplicates (see Fig. 8 in Appendix).
@@ 4.4.1 Semantic versus atomic codes
In Fig. 3 (and Appendix Tab. 6), we report the relative improvement of semantically-structured codes (ger-ald) compared to unstructured codes (ger-atomic). We vary pretraining data size, model capacity and label-space size. A relative improvement of $100\%$ means that the performance of ger-ald doubles compared to ger-atomic.
@@
Entities with long entity names. In Fig. 4 (left), we report the accuracy per entity name length for both ger-ald and ger-caption finetuned models. We see that the longer the entity name, the more ger-ald improves over captioning. Longer entities tend to have more noise with key information further into the code. We also show in Fig. 4 qualitative examples of entities with long entity names (more in Fig. 12 in Appendix). In the left example, we see that ger-ald use the token combination [col][ob] to represent the semantic concept of colobus monkey species. The last token is used to efficiently differentiate between sub-species of colobus. This compact and discriminative way of encoding the entity allows ger-ald to successfully predict this entity whereas ger-caption fails to generate the entity tokenized name.
@@ 4.4.3 Creating codes with ald
Least frequent tokens. In Tab. 3 (left), we validate our choice of selecting the least frequent tokens by evaluating 3 alternatives: random choice, most frequent tokens and first-appearing tokens in tokenized entity name. We see that these alternative strategies hurt the performance significantly. Qualitative examples in Appendix Fig. 11 show that the kept tokens are less semantic and discriminative compared to ger-ald strategy of keeping the least frequent tokens. Note that all these variants are at least as good as ger-atomic ($11.4$ top-1) which is not based on language at all.
@@ 4.5 Link with classification
A typical way of tackling visual entity recognition is by training a classifier into the number of entities [35]. This is not a viable solution for web-scale problems such as OVEN where a single fully-connected layer for a $6$M classes has an enormous parameter count of $4.6$B. In this section, we evaluate ger in cases where learning a classification model is a feasible choice (smaller number of classes). Classification can be cast in our ger framework simply by setting $L=1$ and $V=|\mathcal{E}|=$ number of classes (see Sec. 3.1), making it a special case of atomic codes with $L=1$. Since the decoder decodes a single token, it is equivalent to a multi-layer Multihead Attention Pooling (MAP) head [48, 21]. In Tab. 4, we consider two challenging classification datasets: long-tailed ImageNet-LT [24] and noisy Webvision [22]. We evaluate ger-{ald, atomic} and a classification baseline using multi-layer perceptron (MLP) on averaged-pooled patch tokens. Implementation details are in Sec 6.3 in Appendix.